Mixing 25.0 mL of 1.2 M HCland 25.0 mL of 1.1 M NaOH were mixed. The temperature of the initial solution was 22.4 oC. Assuming a Heat of Neutralization of -55.8 KJ/mol, what would the final temperature be if the specific heat for this solution is 4.03 J/g?
1
Expert's answer
2020-03-04T08:00:24-0500

Q=mc∆T ∆T=Q/mc = -55.8/1.60875 x 4.03 = -8.60677784.

T1-T2=-8.60677784

T2= 22.4 -(-8.60677784) =31.0067778